15-34 Support-department cost allocations; single-department cost pools; direct, step-down, and reciprocal methods. Sportz, Inc., manufactures athletic shoes and athletic clothing for both amateur and professional athletes. The company has two product lines (clothing and shoes), which are produced in separate manufacturing facilities; however, both manufacturing facilities share the same support services for information technology and human resources. The following shows total costs for each manufacturing facility and for each support department. Information technology (IT) Human resources (HR) Clothing Shoes Total costs Variable Costs S 600 S 400 S2,500 $3,000 S6,500 Fixed Costs S 2,000 1,000 S 8,000 S 4,500 $15,500 Total Costs by Department (in thousands) $ 2,600 S 1,400 $10,500 S 7,500 $22,000 The total costs of the support departments (IT and HR) are allocated to the production departments (cloth- ing and shoes) using a single rate based on the following: Information technology Number of IT labor-hours worked by department Number of employees supported by department Human resources. Data on the bases, by department, are given as follows: Department Clothing Shoes Information technology Human resources IT Hours Used 5,040 3,960 Number of Employees 220 92 3,000
1. What are the total costs of the production departments (clothing and shoes) after the support depart- ment costs of information technology and human resources have been allocated using (a) the direct method, (b) the step-down method (allocate information technology first), (c) the step-down method (allocate human resources first), and (d) the reciprocal method? Assume that all of the work of the IT department could be outsourced to an independent company for $97.50 per hour. If Sportz no longer operated its own IT department, 30% of the fixed costs of the IT department could be eliminated. Should Sportz outsource its IT services? 2.
Answer #1
a Direct method
IT department cost 2600
IT hours used(production department only) (5040+3960)
Cost allocation
Cost allocated to Clothing=2600*5040/9000 1456
Cost allocated to Shoes=2600*3960/9000 1144
HR department cost 1400
IT hours used(production department only) (220+88)
Cost allocation
Cost allocated to Clothing=1400*220/308 1000
Cost allocated to Shoes=2600*3960/9000 400
B Step down method(IT Dept first)
Clothing Shoes IT HR
Cost    10,500.00        7,500.00     2,600.00     1,400.00
Allocation of IT department(X/12000*A)      1,092.00            858.00 (2,600.00)         650.00
Total    11,592.00        8,358.00                  -       2,050.00
Allocation of HR department(X/308*A)      1,464.29            585.71 (2,050.00)
Total    13,056.29        8,943.71                  -                    -  
C Step down method(HR Dept first)
Clothing Shoes IT HR
Cost    10,500.00        7,500.00     2,600.00     1,400.00
Allocation of HR department(X/400*A)          770.00            308.00         322.00 (1,400.00)
Total    11,270.00        7,808.00     2,922.00                  -  
Allocation of IT department(X/9000*A)      1,636.32        1,285.68 (2,922.00)
Total    12,906.32        9,093.68                  -                    -  

2.

2 IT department total hour 12000
Current cost 2600
Fixed cost to be reduced 600
Variable cost to be eliminated 600
Unavoidable cost 1400
Cost of outsourcing(in thousands) 1170
Total cost 2570
Cost saving 30

Sports can outsource its IT department as there is a cost saving of 30,000
Answer #2

can anyone solve section D of this problem?? It is the reciprocal method??? It was not included in the answer.

answered by: Anonymous Anonymous